OpenStudy (anonymous):

Solve for x. x/7 + 9 = 15 x = 42 x = 84 x = 168

OpenStudy (alexandervonhumboldt2):

are you joking? x/7=(15-9) x=(15-9)*7

OpenStudy (anonymous):

42?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

thnx